The Beekman Company produces engine parts for car manufacturers. A new accountant intern at Beekman has accidentally deleted the calculations on the​ company's variance analysis calculations for the year ended December​ 31, 2017. The following table is what remains of the data.

Performance Report

Year Ended December 31, 2017

Actual

Flexible-Budget

Flexible

Sales-Volume

Static

Results

Variances

Budget

Variances

Budget

Units sold

103,000

97,000

Revenues (sales)

$695,250

$412,250

Variable costs

450,000

218,250

Contribution margin

245,250

194,000

Fixed costs

156,250

105,000

Operating income

$89,000

$89,000

1.

Calculate all the required variances.​ (If your work is​ accurate, you will find that the total​ static-budget variance is​ $0.)

2.

What are the actual and budgeted selling​ prices? What are the actual and budgeted variable costs per​ unit?

3.

Review the variances you have calculated and discuss possible causes and potential problems. What is the important lesson learned​ her
Answer #1
Performance report
Actual results flexible budget variance flexible budget Sales volume variance static budget
Units sold 103000 0 103000 6000F 97000
Revenues 695250 257500 F 437750 25500F 412250
Variable costs 450000 218250U 231750 13500U 218250
Contribution margin 245250 39250F 206000 12000F 194000
Fixed costs 156250 51250U 105000 0 105000
Operating income/(loss) 89000 12000U 101000 12000F 89000

2.From given static budget

Budgeted selling price = 412250/97000 = 4.25

Budgeted variable cost per unit = 218250/97000 = 2.25

From given actual budget

Actual selling price = 695250/ 103000 = 6.75

Actual variable cost per unit = 450000/103000 = 4.37

3. Flexible budget variance = 12000 U

Sales volume variance = 12000F

Total budget variance = 0